Four years ago, the sum of the ages of A & B was 45. Six years from now, twice B's age will be 28 years more than A's age then. How old is A now?
Answer by Nate(3500)

Four years ago, the sum of the ages of A & B was 45.
A - 4 and B - 4 = 45
A + B - 8 = 45
A + B = 53
Six years from now, twice B's age will be 28 years more than A's age then. How old is A now?
2(B + 6) = 28 + (A + 6)
2B + 12 = 34 + A
2B - 22 = A
Plug:
A + B = 53
2B - 22 + B = 53
3B = 75
B = 25
Plug:
A + B = 53
A + 25 = 53
A = 28